OpenStudy (anonymous):

find the volume of the solid generated by the revolution of the curve y^2 (2 a - x) = x^3 about its asymptote.

HanAkoSolo (jamierox4ev3r):

@austin19 @AddingValue @Ambbiiee @blurbendy @bngaure @bernadith @ChanceBBrick @Danny_Boy @druminjosh @ess76 @ericmyer @feryy_ @Darion1234 @whpalmer4 @radar @t3l4ibfw4in49g8g435g4gg5g

OpenStudy (kainui):

How far have you gotten? Can you find the asymptote? We'll take it step by step.

OpenStudy (anonymous):

i tried using disk method. but its not not working because of equations format

OpenStudy (kainui):

Sure, show me what equation you came up with to integrate with the disk method. Make sure you solve for y before you try disk method.

OpenStudy (anonymous):

you mean generalized equation ?

OpenStudy (kainui):

I think so, not sure what you mean by generalized equation. Just try to get it looking like a formula for y= something in terms of x.

OpenStudy (anonymous):

y=x^(3/2)/[2a-x]^(1/2)

OpenStudy (dumbcow):

i would use shell method here so you can integrate wrt "x" its still a messy integral

OpenStudy (dumbcow):

\[V = 4\pi \int\limits _{0} ^{2a} (2a-x)\frac{x^{3/2}}{\sqrt{2a-x}} dx\]

OpenStudy (anonymous):

how did you get the limit 0 to 2a ?

OpenStudy (dumbcow):

vertical asymptote is x=2a domain of x varies from 0 to 2a

OpenStudy (anonymous):

and 2a-x is the radius of the shell right ?

OpenStudy (dumbcow):

correct and height is "y"

OpenStudy (anonymous):

thank you.

OpenStudy (dumbcow):

yw, also due to symmetry you only integrate for half volume then double it (thtas where the 4 came from)

OpenStudy (anonymous):

can you suggest any study material in shell method.